Dimensional Analysis and Units


While driving in the U.S.A., a Canadian fuels their car at a cost of US$3.08 per gallon (US).
Given that the monetary exchange rate is 0.693 US$/Can$, and the density of the fuel is 0.626 g/mL, what is the mass of Can$50.00-worth of this fuel?

1pts

Tries 0/5

If the vehicle consumes the gasoline at a rate of 5.96 L/100 km, how many miles can it travel on the above amount of gasoline?
miles
Answer #1

Problem 1.

1 gallon = 3785.41 mL

3.08 US$ = 3.08/0.693 Can$ = 4.444 Can$

The fuel for 50 Can$ = (50/4.444)* 3785.41 mL = 42590 mL = 42.59 L

Density of fuel = 0.626 g/mL

Therefore, the required mass of fuel = 42590 mL * 0.626 g/mL = 26661.34 g = 26.66 kg

Problem 2.

1 mile = 1.60934 km

The rate of gasoline consumption = 5.96 L/62.1371 mile

Therefore, the required no. of miles = 42.59 L/(5.96 L/62.1371 mile) ~ 444 miles